# FedMentor: Domain-Aware Differential Privacy for Heterogeneous Federated LLMs in Mental Health

### v2

@@ 1 Introduction
There is growing interest across the AI and healthcare communities in developing trustworthy and scalable mental health chatbots that can provide rapid, accessible, and confidential psychological support. In 2019, mental disorders affected about 970 million people worldwide, or one in eight individuals Institute for Health Metrics and Evaluation (2023) (IHME). Conversational agents powered by large language models have since emerged as promising tools for supporting mental well-being at scale Towards Healthcare (2023). The global market for mental health chatbots, valued at \$0.99 billion in 2022, is projected to grow to \$6.51 billion by 2032, underscoring strong clinical and commercial demand Towards Healthcare (2023). At the same time, deploying LLMs in this setting raises critical challenges in privacy protection, legal compliance, and efficient communication and computation for personalized care Lai et al. (2023); United States Congress (1996); European Parliament and Council of the European Union (2016). These challenges have motivated research into privacy-preserving learning, parameter-efficient adaptation, and federated methods tailored for high-risk domains such as mental health, where user trust and confidentiality are essential.
@@
To investigate privacy and communication challenges in adapting LLMs to sensitive mental health domains, we present FedMentor, a federated framework that assigns domain-aware Differential Privacy budgets and fine-tunes only LoRA adapters, aggregated with FedAvg McMahan et al. (2017), to achieve private adaptation while preserving utility and fairness under heterogeneity. The overall architecture is shown in Figure 2. FedMentor is designed for scenarios that demand distinct privacy guarantees across domains and where non-IID data can introduce drift and unequal outcomes. We evaluate the framework on three datasets (Dreaddit Turcan and McKeown (2019), IRF Garg et al. (2023), MultiWD Garg et al. (2024)) and five backbones (MobileLLM-ParetoQ-350M Liu et al. (2024b), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025)). FedMentor improves safety with minimal loss in utility: for instance, on MultiWD the Toxicity Safe Rate (TSR) rises by two points while toxicity decreases from 2.92 to 1.98 (Table 1), and BERTScore F1 Zhang et al. (2020) and ROUGE-L Lin (2004) remain close to standard FL and near the centralized upper bound. Client-level relevance remains consistent across domains, and efficiency is achieved by communicating only adapter weights through LoRA, keeping computation and bandwidth feasible for single-GPU clients. To the best of our knowledge, this is the first study that combines domain-aware DP Dwork et al. (2014); Abadi et al. (2016) with federated LoRA fine-tuning for large language models in mental health.
@@ 4 Experiment Settings
Datasets. We evaluate FedMentor on three mental health datasets: Dreaddit Turcan and McKeown (2019) for stress detection, Interpersonal Risk Factors (IRF) Garg et al. (2023) for Thwarted Belongingness (TBe) and Perceived Burdensomeness (PBu), and MultiWD Garg et al. (2024) for multi-label wellness prediction. Together, these corpora cover complementary tasks and form a natural non-IID benchmark across domains. Regulatory requirements under HIPAA and GDPR restrict centralized collection, and lengthy institutional approvals further limit access, producing small, fragmented datasets Act et al. (1996); Regulation (2018); Nicholas et al. (2020). Prior FL work mitigates performance degradation from heterogeneity with methods such as SCAFFOLD Karimireddy et al. (2020), FedNova Wang et al. (2020), and Adaptive FedOpt Reddi et al. (2021). We adopt a domain-aware FL setting where each dataset defines a client domain and use this setting to evaluate FedMentor under heterogeneous conditions. App. B.1 provides dataset statistics and preprocessing details.
@@ 5 Main results
Baseline
Domain
Spec.
2.0 / 0.5 / 1.5 derived $97.33$ $94$ $100$ $2.49$ 66
IRF Sensitivity
Domain
Spec.
fixed 2.0 / 0.1 / 1.5 derived $78.33$ $66$ $88$ $9.18$ $22$
@@
Utility Threshold Global
B-F1
$\tau{=}0$
$74.67$ $65$ $85$ $8.18$ 2020
@@ 6 Ablation Studies
Uniform Privacy. Uniform Privacy applies a single global privacy budget to all domains. In Figure 4a–b and Table 3, this choice markedly improves fairness in safety: with a uniform $\epsilon{=}1.0$, the TSR concentrates at $\sim$99–100% and the client spread contracts to 1% (vs. 6% under the domain-specific baseline). In contrast, the static noise variant shows lower average safety (87.33% TSR) and a larger spread (18%). The utility-gated variant with $\tau{=}0$ yields the lowest safety (74.67% TSR). BERTScore F1 and REL remain close to the baseline across settings, indicating that the fairness gains from a uniform budget come with only modest utility change. All panels in Figure 4 report the final global model after 8 rounds.
@@
Per-Client Memory Scaling across Rounds Table 5 reports memory usage per client for Qwen3-1.7B under the IRF $\epsilon{=}1.0$ setting. Memory increases smoothly over rounds: at round 0, clients use about 18.85, 20.04, and 21.23 GB; by round 7, these reach 43.78, 44.97, and 46.16 GB. The monotonic rise indicates predictable scaling during training while keeping per-client usage within a narrow band across clients.
7 Related Works
@@ Appendix A Additional Related Works
Differential Privacy in FL. Differential privacy (DP) has become a standard approach for protecting sensitive data in ML, including FL Dwork et al. (2014); Abadi et al. (2016); Girgis et al. (2021); Noble et al. (2022). Applying DP in practice, however, remains difficult, particularly in mental health settings. Traditional methods that uniformly clip gradients and add noise to updates often degrade performance in healthcare, where datasets are small and signals are fine-grained McMahan et al. (2018). Although DP has been studied in areas such as electronic health records and medical imaging, its application in mental health remains limited, with only a few recent studies exploring privacy-by-design strategies Choudhury et al. (2019); Sawhney et al. (2023); Romijnders et al. (2024). Recent research further highlights that not all model components or data are equally sensitive, motivating personalized and domain-aware DP strategies that adapt privacy to data characteristics Ebadi et al. (2015); Jorgensen et al. (2015); Liu et al. (2021, 2022, 2024a). In LLMs, parameter-efficient methods such as DP-LoRA Liu et al. (2025) reduce risk and communication costs by perturbing only low-rank adapter updates. However, most approaches still adopt a one-size-fits-all noise and overlook the unique sensitivities of mental health text. To overcome this limitation, we introduce FedMentor, a framework that integrates domain-aware DP with heterogeneous federated learning to provide strong privacy guarantees while preserving clinical utility.
Appendix B Additional Experiment Settings
B.1 Dataset Details
(a) Dreaddit
(b) IRF
(c) MultiWD
B.2 Model Specification
B.3 Baseline Specification

### v3

@@ 1 Introduction
There is growing interest across the AI and healthcare communities in developing trustworthy and scalable mental health chatbots that can provide rapid, accessible, and confidential psychological support. In 2019, mental disorders affected about 970 million people worldwide, or one in eight individuals Institute for Health Metrics and Evaluation (IHME) (2023). Conversational agents powered by large language models have since emerged as promising tools for supporting mental well-being at scale Towards Healthcare (2023). The global market for mental health chatbots, valued at \$0.99 billion in 2022, is projected to grow to \$6.51 billion by 2032, underscoring strong clinical and commercial demand Towards Healthcare (2023). At the same time, deploying LLMs in this setting raises critical challenges in privacy protection, legal compliance, and efficient communication and computation for personalized care Lai et al. (2023); United States Congress (1996); European Parliament and Council of the European Union (2016). These challenges have motivated research into privacy-preserving learning, parameter-efficient adaptation, and federated methods tailored for high-risk domains such as mental health, where user trust and confidentiality are essential.
@@
To investigate privacy and communication challenges in adapting LLMs to sensitive mental health domains, we present FedMentor, a federated framework that assigns domain-aware Differential Privacy budgets and fine-tunes only LoRA adapters, aggregated with FedAvg McMahan et al. (2017), to achieve private adaptation while preserving utility and fairness under heterogeneity. The overall architecture is shown in Figure 2. FedMentor is designed for scenarios that demand distinct privacy guarantees across domains and where non-IID data can introduce drift and unequal outcomes. We evaluate the framework on three datasets (Dreaddit Turcan and McKeown (2019), IRF Garg et al. (2023), MultiWD Garg et al. (2024)) and five backbones (MobileLLM-ParetoQ-350M Liu et al. (2024b), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen3 (0.6B and 1.7B) Yang et al. (2025)). FedMentor improves safety with minimal loss in utility: for instance, on MultiWD the Toxicity Safe Rate (TSR) rises by two points while toxicity decreases from 2.92 to 1.98 (Table 1), and BERTScore F1 Zhang et al. (2020) and ROUGE-L Lin (2004) remain close to standard FL and near the centralized upper bound. Client-level relevance remains consistent across domains, and efficiency is achieved by communicating only adapter weights through LoRA, keeping computation and bandwidth feasible for single-GPU clients. To the best of our knowledge, this is the first study that combines domain-aware DP Dwork et al. (2014); Abadi et al. (2016) with federated LoRA fine-tuning for LLMs in mental health.
@@ 4 Experiment Settings
Datasets. We evaluate FedMentor on three mental health datasets: Dreaddit Turcan and McKeown (2019) for stress detection, Interpersonal Risk Factors (IRF) Garg et al. (2023) for Thwarted Belongingness (TBe) and Perceived Burdensomeness (PBu), and MultiWD Garg et al. (2024) for multi-label wellness prediction. Together, these corpora cover complementary tasks and form a natural non-IID benchmark across domains. Regulatory requirements under HIPAA and GDPR restrict centralized collection, and lengthy institutional approvals further limit access, producing small, fragmented datasets Act and others (1996); Regulation (2018); Nicholas et al. (2020). Prior FL work mitigates performance degradation from heterogeneity with methods such as SCAFFOLD Karimireddy et al. (2020), FedNova Wang et al. (2020), and Adaptive FedOpt Reddi et al. (2021). We adopt a domain-aware FL setting where each dataset defines a client domain and use this setting to evaluate FedMentor under heterogeneous conditions. App. B.1 provides dataset statistics and preprocessing details.
@@ 5 Main results
Baseline Domain Spec. 2.0 / 0.5 / 1.5 derived $97.33$ $94$ $100$ $2.49$ 66
IRF Sensitivity Domain Spec. fixed 2.0 / 0.1 / 1.5 derived $78.33$ $66$ $88$ $9.18$ $22$
@@
Utility Threshold Global B-F1 $\tau{=}0$ $74.67$ $65$ $85$ $8.18$ 2020
@@ 6 Ablation Studies
Uniform Privacy. Uniform Privacy applies a single global privacy budget to all domains. In Figure 4a-b and Table 3, this choice markedly improves fairness in safety: with a uniform $\epsilon{=}1.0$, the TSR concentrates at $\sim$99-100% and the client spread contracts to 1% (vs. 6% under the domain-specific baseline). In contrast, the static noise variant shows lower average safety (87.33% TSR) and a larger spread (18%). The utility-gated variant with $\tau{=}0$ yields the lowest safety (74.67% TSR). BERTScore F1 and REL remain close to the baseline across settings, indicating that the fairness gains from a uniform budget come with only modest utility change. All panels in Figure 4 report the final global model after 8 rounds.
@@
Per-Client Memory Scaling across Rounds. Table 5 reports memory usage per client for Qwen3-1.7B under the IRF $\epsilon{=}1.0$ setting. Memory increases smoothly over rounds: at round 0, clients use about 18.85, 20.04, and 21.23 GB; by round 7, these reach 43.78, 44.97, and 46.16 GB. The monotonic rise indicates predictable scaling during training while keeping per-client usage within a narrow band across clients.
@@ Appendix A Additional Related Works
Differential Privacy in FL. Differential privacy (DP) has become a standard approach for protecting sensitive data in ML, including FL Dwork et al. (2014); Abadi et al. (2016); Girgis et al. (2021); Noble et al. (2022). Applying DP in practice, however, remains difficult, particularly in mental health settings. Traditional methods that uniformly clip gradients and add noise to updates often degrade performance in healthcare, where datasets are small and signals are fine-grained McMahan et al. (2018). Although DP has been studied in areas such as electronic health records and medical imaging, its application in mental health remains limited, with only a few recent studies exploring privacy-by-design strategies Choudhury et al. (2019); Sawhney et al. (2023); Romijnders et al. (2024). Recent research further highlights that not all model components or data are equally sensitive, motivating personalized and domain-aware DP strategies that adapt privacy to data characteristics Ebadi et al. (2015); Jorgensen et al. (2015); Liu et al. (2021, 2022, 2024a). In LLMs, parameter-efficient methods such as DP-LoRA Liu et al. (2025) reduce risk and communication costs by perturbing only low-rank adapter updates. However, most approaches still adopt a one-size-fits-all noise and overlook the unique sensitivities of mental health text. To overcome this limitation, we introduce FedMentor, a framework that integrates domain-aware DP with heterogeneous FL to provide strong privacy guarantees while preserving clinical utility.
@@
Appendix B Additional Experiment Settings
B.1 Dataset Details
(a) Dreaddit
(b) IRF
(c) MultiWD
B.2 Model Specification
B.3 Baseline Specification